# MAGE: Model-Level Graph Neural Networks Explanations via Motif-based Graph Generation

### v1

@@ Abstract
Graph Neural Networks (GNNs) have shown remarkable success in molecular tasks, yet their interpretability remains challenging. Traditional model-level explanation methods like XGNN and GNNInterpreter often fail to identify valid substructures like rings, leading to questionable interpretability. This limitation stems from XGNN’s atom-by-atom approach and GNNInterpreter’s reliance on average graph embeddings, which overlook the essential structural elements crucial for molecules. To address these gaps, we introduce an innovative Motif-bAsed GNN Explainer (MAGE) that uses motifs as fundamental units for generating explanations. Our approach begins with extracting potential motifs through a motif decomposition technique. Then, we utilize an attention-based learning method to identify class-specific motifs. Finally, we employ a motif-based graph generator for each class to create molecular graph explanations based on these class-specific motifs. This novel method not only incorporates critical substructures into the explanations but also guarantees their validity, yielding results that are human-understandable. Our proposed method’s effectiveness is demonstrated through quantitative and qualitative assessments conducted on six real-world molecular datasets.
@@ 1 Introduction
Graph Neural Networks (GNNs) [18, 39, 12, 28, 33, 38] have become increasingly popular tools for modeling data in the molecule field. As an effective method for learning representations from molecule data, GNNs have attained state-of-the-art results in tasks like molecular representation learning [11, 43, 10, 48], and molecule generation [6, 20]. Despite their growing popularity, questions arise about the trustworthiness and decision-making processes of GNNs.
Explaining GNNs has become a major area of interest in recent years and existing methods can be broadly categorized into instance-level explanations and model-level explanations. Instance-level explanations [42, 4, 26, 13, 34, 29, 22, 49, 30, 36, 47, 35, 44] aim to pinpoint specific nodes, edges, or subgraphs crucial for a GNN model’s predictions on one data instance. Model-level explanations [46, 37] seek to demystify the overall behavior of the GNN model by identifying patterns that generally lead to certain predictions. While instance-level methods offer detailed insights, they require extensive analysis across numerous examples to be deemed reliable. Conversely, model-level methods provide more high-level explanations that require less intensive human oversight.
Model-level explanation methods have two main categories: concept-based methods and generation-based methods. Concept-based explanation methods focus on identifying higher-level concepts that significantly influence the model’s predictions [3, 40]. Also, these methods establish rules to illustrate how these concepts are interconnected, like using logical formulas [3]. Generation-based models try to learn a generative model that can generate synthetic graphs that are optimized to maximize the behavior of the target GNN [46, 37]. Instead of defining the relationships between important concepts with formulas, generation-based models can generate novel graph structures that are optimized for specific properties. This capability is particularly crucial in the field of molecular learning.
@@
Despite the notable benefits of generation-based model-level explanation methods for GNNs, they have received relatively low attention. Presently, the available generation-based model-level explainers for GNNs are somewhat limited in their application, especially in molecular graphs. For example, XGNN [46] explains models by building an explanation atom-by-atom and sets a maximum degree for each atom to maintain the explanation’s validity. On the other hand, GNNInterpreter [37] adopts a more flexible approach by learning a generative explanation graph distribution. However, it ensures the validity of explanations by maximizing the similarity between the explanation graph embedding and the average embedding of all graphs, which is not particularly effective for molecular graphs. These existing methods often fail to consider the unique aspects of molecular structures adequately. This oversight makes it challenging for these methods to include crucial substructures in their explanation graphs. It compromises the validity and reliability of the model-level explanations they provide for molecular graphs. In Figure 1, the explanations produced by both XGNN and GNNInterpreter exhibit low validity, making them ineffective for further analysis. Consequently, there’s a clear need for more advanced and specialized model-level explanation methods that can effectively handle the unique characteristics of molecular graphs and provide meaningful insights into how GNNs interpret and process molecular data.
@@ 2 Preliminary
Notations. A molecular dataset can be depicted as $\mathcal{G}=\{G_{1},G_{2},...,G_{i},...,G_{n}\}$, where each molecular graph within this set is denoted by $G_{i}=(\textbf{V},\textbf{E})$. The label of each graph will be assigned by $l_{i}\in L=\{l_{1},l_{2},...\}$ In this representation, V constitutes a set of atoms, each treated as a node, while E comprises the bonds, represented as edges connecting these nodes. The structural information of each graph $G_{i}$ is encoded in several matrices. The adjacency matrix $\textbf{A}\in\{0,1\}^{N\times N}$ represents the connections between atoms, where a ‘1’ indicates the presence of a bond between a pair of atoms, a ‘0’ signifies no bond, and $N$ represents the total number of atoms in the graph. Additionally, an atom feature matrix $\textbf{X}\in\mathbb{R}^{N\times D}$ captures the features of each atom, with $D$ being the number of features characterizing each atom. Similarly, a bond feature matrix $\textbf{Z}\in\mathbb{R}^{N\times D_{b}}$ is used to describe the attributes of the bonds, where $D_{b}$ denotes the number of features describing each bond.
Graph Neural Networks. Graph Neural Networks (GNNs) leverage adjacency, node feature, and edge feature matrices to learn node features. Despite the existence of various GNN variants like Graph Convolutional Networks (GCNs), Graph Attention Networks (GATs), and Graph Isomorphism Networks (GINs), they generally follow a message-passing mechanism. In this process, each GNN layer aggregates information from a node’s neighbors to refine its representation. For a given hidden layer $i$, the message passing operation can be formulated as $h_{i}=f(h_{i-1},A,Z)$, where $h_{0}=X$. Here, $h_{i}\in\mathbb{R}^{N\times D}$ is the node representation output from layer $i$, and $D$ is the dimension of the output features. The function $f$ involves computing messages based on the previous layer’s node embeddings ($h_{i-1}$), aggregating these messages from neighboring nodes, and updating the hidden representation $h_{i}$ for each node using the aggregated messages.
Motif. A motif is essentially a recurring subgraph within a graph, closely associated with the graph’s property, which has been thoroughly researched across various domains, including biochemistry, ecology, neurobiology, and engineering [24, 31, 1, 2] and has been demonstrated to be of significant importance. GSN [7] introduces motif information into node features to improve the learning of GNNs. HMGNN [43] generates a heterogeneous motif graph to enhance the molecular representation learning. MotifExplainer [44] applies motifs to learn an instance-level explanation for GNNs.
@@ 3 Motif-based Model-Level GNN Explainer
Given a dataset, denoted as $\mathcal{G}$ with $N$ molecules and $C$ classes, our objective is to generate model-level explanations for a target GNN, which comprises a feature extractor $\phi(\cdot)$ and a classifier $f(\cdot)$. $\phi(\cdot)$ takes a molecular graph or subgraph as input and outputs its feature representation. $f(\cdot)$ utilizes this representation to predict the classification of a molecule.
@@ 3.1 Motif Extraction
There are mainly three methods for motif extraction. It’s important to note that these motif extraction methods can be integrated into our MAGE explanation method, demonstrating its versatility.
1) Ring and bonded pairs (R&B). Methods [15, 43, 7] in this category recognize simple rings and bonded pairs as motifs. 2) Molecule decomposition. In this category, methods recognize motifs through molecule breakdown. These methods initially establish a catalog of target bonds, either through domain-specific knowledge, such as the “breakable bonds” used by RECAP [21, 50] and BRICS [8, 48, 14], or through hand-crafted rules, like “bridge bonds” [16]. The algorithms remove these target bonds from the molecules and use the consequent molecular fragments as motifs. 3) Molecule tokenization. Certain methods [9, 19] leverage the string representations of molecules, such as SMILES, and directly apply the WordPiece algorithm from the NLP domain. The resulting strings within the constructed vocabulary are regarded as motifs. 4) Data-driven method. MotifPiece [45] statistically identifies underlying structural or functional patterns specific to a given molecular data.
@@ 3.2 Class-wise Motif Identification
$\boldsymbol{h}_{G_{i}}=\phi(G_{i}),\boldsymbol{h}_{m_{i}}=\phi(m_{i}),\text{ for all $m_{i}$}\in M_{i}.$
Next, an attention operator is used to aggregate the motif encoding using $\boldsymbol{h}_{G_{i}}$ as query and $\boldsymbol{h}_{m_{i}}$s as keys and values. The resulting encoding, denoted as $\boldsymbol{h}^{\prime}_{G_{i}}$, is the aggregated molecular encoding by combining motifs in $M_{i}$.
$e_{G_{i}m_{i}}=g(\boldsymbol{W}\boldsymbol{h}_{G_{i}},\boldsymbol{W}\boldsymbol{h}_{m_{i}}),\alpha_{G_{i}m_{i}}=\frac{\exp(e_{G_{i}m_{i}})}{\sum_{m_{i}\in M_{i}}\exp(e_{G_{i}m_{i}})},\boldsymbol{h}^{\prime}_{G_{i}}=\sum_{m_{i}\in M_{i}}\alpha_{G_{i}m_{i}}\cdot\boldsymbol{h}_{m_{i}},$
where $g(\cdot)$ is an attention operator, and $\boldsymbol{W}$ is weight matrix. The target classifier $f(\cdot)$ then uses aggregated encoding to generate predictions. The training process is designed to minimize the cross-entropy loss between the predicted probabilities of the aggregated molecular encoding and those of the original molecular encoding:
$\mathcal{L}=\sum_{i}\mathcal{L}(f(\boldsymbol{h}_{G_{i}}),f(\boldsymbol{h}^{\prime}_{G_{i}})).$ (1)
The attention operator is trained to identify important motifs to represent the molecule. This leads to the molecule-motif relation score $\boldsymbol{\alpha}_{G_{i}m_{i}}$ and the probabilities associated with label predictions $\boldsymbol{\hat{y}}=f(\boldsymbol{h}^{\prime}_{G_{i}})$.
@@
We repeat the process of Stage 1 to compute the molecule-motif scores for each molecular graph. After that, we create a molecule-motif score matrix $\boldsymbol{S}^{mm}\in\mathbb{R}^{V\times N}$ and a molecule-class probability matrix $\boldsymbol{P}\in\mathbb{R}^{N\times C}$. Here, ${S}^{mm}_{pk}$ represents the molecule-motif relation score between molecule $k$ and motif $p$. If the motif $i$ does not appear in the molecule $j$, we set ${S}^{mm}_{ij}=0$. Element ${P}_{kr}$ denotes the probability that molecule $k$ is associated with label $r$.
Subsequently, we perform a matrix multiplication between $\boldsymbol{S}^{mm}$ and $\boldsymbol{P}$ to get class-motif relation score matrix $\boldsymbol{S}^{cm}$:
$\boldsymbol{S}^{cm}=\boldsymbol{S}^{mm}\boldsymbol{P},{S}^{cm}_{pr}=\sum_{k=1}^{N}{S}^{mm}_{pk}{P}_{kr}$
which represents the relationship between motif $p$ and class $r$. Then, we normalize $\boldsymbol{S}^{cm}$ by dividing each column, corresponding to a motif, by the number of molecules it appears.
@@ 3.3 Class-wise Motif-based Graph Generation
Figure 3: Illustration of class-wise motif-based graph generation.
Section 3.2 identifies motifs of high relevance for each class. Building on this, our methodology employs a motif-based generative model to construct model-level explanations tailored to each class. The motif-based molecular generation method can be divided into tree-constrained [15, 16] and non-constrained methods [41]. In our approach, we employ a tree-constrained method as a generator. However, our approach can easily adapt to non-constrained methods. The graph generation process is illustrated in Figure 3.
@@ 3.3.2 Graph Encoder
We encode the latent representation of $G=(\textbf{A},\textbf{X})$ by the target GNN feature extractor $\phi(\cdot)$: $\boldsymbol{h}_{G}=\phi(\textbf{A},\textbf{X})$. The mean $\boldsymbol{\mu}_{G}$ and log variance $\log\boldsymbol{\sigma}_{G}$ of the variational posterior approximation are computed from $\boldsymbol{h}_{G}$ with two separate affine layers. $\boldsymbol{z}_{G}$ is sampled from a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu}_{G},\boldsymbol{\sigma}_{G})$. In the variational posterior approximation, we use two different affine layers to compute the mean, represented as $\boldsymbol{\mu}_{G}$, and the log variance, denoted as $\log\boldsymbol{\sigma}_{G}$, from $\boldsymbol{h}_{G}$. Then, we sample $\boldsymbol{z}_{G}$ from a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu}_{G},\boldsymbol{\sigma}_{G})$.

### v2

@@ Abstract
Graph Neural Networks (GNNs) have shown remarkable success in molecular tasks, yet their interpretability remains challenging. Traditional model-level explanation methods like XGNN and GNNInterpreter often fail to identify valid substructures like rings, leading to questionable interpretability. This limitation stems from XGNN’s atom-by-atom approach and GNNInterpreter’s reliance on average graph embeddings, which overlook the essential structural elements crucial for molecules. To address these gaps, we introduce an innovative Motif-bAsed GNN Explainer (MAGE) that uses motifs as fundamental units for generating explanations. Our approach begins with extracting potential motifs through a motif decomposition technique. Then, we utilize an attention-based learning method to identify class-specific motifs. Finally, we employ a motif-based graph generator for each class to create molecular graph explanations based on these class-specific motifs. This novel method not only incorporates critical substructures into the explanations but also guarantees their validity, yielding results that are human-understandable. Our proposed method’s effectiveness is demonstrated through quantitative and qualitative assessments conducted on six real-world molecular datasets. The implementation of our method can be found at https://github.com/ZhaoningYu1996/MAGE.
@@ 1 Introduction
Graph Neural Networks (GNNs) (Kipf & Welling, 2016; Xu et al., 2018; Gao et al., 2018; Rong et al., 2020; Sun et al., 2022; Wang et al., 2023) have become increasingly popular tools for modeling data in the molecule field. As an effective method for learning representations from molecule data, GNNs have attained state-of-the-art results in tasks like molecular representation learning (Gao & Ji, 2019; Yu & Gao, 2022a; Fang et al., 2022; Zang et al., 2023), and molecule generation (Bongini et al., 2021; Lai et al., 2021). Despite their growing popularity, questions arise about the trustworthiness and decision-making processes of GNNs.
Explaining GNNs has become a major area of interest in recent years and existing methods can be broadly categorized into instance-level explanations and model-level explanations. Instance-level explanations (Ying et al., 2019; Baldassarre & Azizpour, 2019; Pope et al., 2019; Huang et al., 2020; Vu & Thai, 2020; Schnake et al., 2020; Luo et al., 2020; Funke et al., 2020; Zhang et al., 2021a; Shan et al., 2021; Wang et al., 2021; Yuan et al., 2021; Wang et al., 2022; Yu & Gao, 2022b) aim to pinpoint specific nodes, edges, or subgraphs crucial for a GNN model’s predictions on one data instance. Model-level explanations (Yuan et al., 2020; Wang & Shen, 2022) seek to demystify the overall behavior of the GNN model by identifying patterns that generally lead to certain predictions. While instance-level methods offer detailed insights, they require extensive analysis across numerous examples to be reliable. In addition, for instance-level explainer, there is no requirement for an explanation to be valid (e.g., a chemically valid molecular graph). Conversely, model-level methods provide more high-level and broader insights explanations that require less intensive human oversight.
Model-level explanation methods have two main categories: concept-based methods and generation-based methods. Concept-based explanation methods focus on identifying higher-level concepts that significantly influence the model’s predictions (Azzolin et al., 2022; Xuanyuan et al., 2023). Also, these methods establish rules to illustrate how these concepts are interconnected, like using logical formulas (Azzolin et al., 2022). Generation-based models try to learn a generative model that can generate synthetic graphs that are optimized to maximize the behavior of the target GNN (Yuan et al., 2020; Wang & Shen, 2022). Instead of defining the relationships between important concepts with formulas, generation-based models can generate novel graph structures that are optimized for specific properties. This capability is particularly crucial in the field of molecular learning.
@@
Despite the notable benefits of generation-based model-level explanation methods for GNNs, they have received relatively low attention. Presently, the available generation-based model-level explainers for GNNs are somewhat limited in their application, especially in molecular graphs. For example, XGNN (Yuan et al., 2020) explains models by building an explanation atom-by-atom and sets a maximum degree for each atom to maintain the explanation’s validity. On the other hand, GNNInterpreter (Wang & Shen, 2022) adopts a more flexible approach by learning a generative explanation graph distribution. However, it ensures the validity of explanations by maximizing the similarity between the explanation graph embedding and the average embedding of all graphs, which is not particularly effective for molecular graphs. These existing methods often fail to consider the unique aspects of molecular structures. This oversight makes it challenging for these methods to include crucial substructures in their explanation graphs. It compromises the validity and reliability of the model-level explanations they provide for molecular graphs. In Figure 1, the explanations produced by both XGNN and GNNInterpreter exhibit low validity, making them ineffective for further analysis. Consequently, there’s a clear need for more advanced and specialized model-level explanation methods that can effectively handle the unique characteristics of molecular graphs and provide meaningful insights into how GNNs interpret and process molecular data.
@@
The main contribution of our paper can be summarized as below:
  1. 1.
    We propose a novel framework for model-level explanations of GNNs that produces chemical valid explanations on molecular graphs.
  2. 2.
    We develop an attention-based mechanism to extract class-specific motifs, allowing MAGE to provide tailored model-level explanations aligned with the predictive behavior of the target GNNs for each class.
  3. 3.
    We employ motifs as foundational building blocks for model-level explanations generation, offering interpretable and structurally coherent insights into GNN behavior.
  4. 4.
    We demonstrate MAGE’s superiority over existing methods through quantitative and qualitative evaluations on multiple molecular datasets.
@@ 2 Preliminary
Notations. A molecular dataset can be depicted as $\mathcal{G}=\{G_{1},G_{2},...,G_{i},...,G_{n}\}$, where each molecular graph within this set is denoted by $G_{i}=(V,E)$. The label of each graph will be assigned by $l_{i}\in L=\{l_{1},l_{2},...\}$. In this representation, $V$ constitutes a set of atoms, each treated as a node, while $E$ comprises the bonds, represented as edges connecting these nodes. The structural information of each graph $G_{i}$ is encoded in several matrices. The adjacency matrix $\bm{A}\in\{0,1\}^{N\times N}$ represents the connections between atoms, where a ‘1’ indicates the presence of a bond between a pair of atoms, a ‘0’ signifies no bond, and $N$ represents the total number of atoms in the graph. Additionally, an atom feature matrix $\bm{X}\in\mathbb{R}^{N\times D}$ captures the features of each atom, with $D$ being the number of features characterizing each atom. Similarly, a bond feature matrix $\bm{Z}\in\mathbb{R}^{|E|\times D_{b}}$ is used to describe the attributes of the bonds, where $|E|$ represents the number of edges and $D_{b}$ denotes the number of features describing each bond.
Graph Neural Networks. Graph Neural Networks (GNNs) leverage adjacency matrix, node feature matrix, and edge feature matrix to learn node features. Despite the existence of various GNN variants like Graph Convolutional Networks (GCNs), Graph Attention Networks (GATs), and Graph Isomorphism Networks (GINs), they generally follow a message-passing mechanism. In this process, each GNN layer aggregates information from a node’s neighbors to refine its representation. For a given hidden layer $i$, the message passing operation can be formulated as $h_{i}=f(h_{i-1},\bm{A},\bm{Z})$, where $h_{0}=X$. Here, $h_{i}\in\mathbb{R}^{N\times D}$ is the node representation output from layer $i$, and $D$ is the dimension of the output features. The function $f$ involves computing messages based on the previous layer’s node embeddings ($h_{i-1}$), aggregating these messages from neighboring nodes, and updating the hidden representation $h_{i}$ for each node using the aggregated messages.
Motif. A motif is essentially a recurring subgraph within a graph, closely associated with the graph’s property, which has been thoroughly researched across various domains, including biochemistry, ecology, neurobiology, and engineering (Milo et al., 2002; Shen-Orr et al., 2002; Alon, 2007; 2019) and has been demonstrated to be of significant importance. GSN (Bouritsas et al., 2022) introduces motif information into node features to improve the learning of GNNs. HMGNN (Yu & Gao, 2022a) generates a heterogeneous motif graph to enhance the molecular representation learning. MotifExplainer (Yu & Gao, 2022b) applies motifs to learn an instance-level explanation for GNNs.
@@ 3 Motif-based Model-Level GNN Explainer
Given a dataset, denoted as $\mathcal{G}$ with $|\mathcal{G}|$ molecules and $C$ classes, our objective is to generate model-level explanations for a target GNN, which comprises a feature extractor $\phi(\cdot)$ and a classifier $f(\cdot)$. $\phi(\cdot)$ takes a molecular graph or subgraph as input and outputs its feature representation. $f(\cdot)$ utilizes this representation to predict the classification of a molecule.
@@ 3.1 Motif Extraction
There are mainly four methods for motif extraction. It’s important to note that these motif extraction methods can be integrated into our MAGE explanation method, demonstrating its versatility.
1) Ring and bonded pairs (R&B). Methods (Jin et al., 2018; Yu & Gao, 2022a; Bouritsas et al., 2022) in this category recognize simple rings and bonded pairs as motifs. 2) Molecule decomposition. In this category, methods recognize motifs through molecule breakdown. These methods initially establish a catalog of target bonds, either through domain-specific knowledge, such as the “breakable bonds” used by RECAP (Lewell et al., 1998; Zhang et al., 2021b) and BRICS (Degen et al., 2008; Zang et al., 2023; Jiang et al., 2023), or through hand-crafted rules, like “bridge bonds” (Jin et al., 2020). The algorithms remove these target bonds from the molecules and use the consequent molecular fragments as motifs. 3) Molecule tokenization. Certain methods (Fang et al., 2023; Kuenneth & Ramprasad, 2023) leverage the string representations of molecules, such as SMILES, and directly apply the WordPiece algorithm from the NLP domain. The resulting strings within the constructed vocabulary are regarded as motifs. 4) Data-driven method. MotifPiece (Yu & Gao, 2023) statistically identifies underlying structural or functional patterns specific to a given molecular data.
@@ 3.2 Class-wise Motif Identification
$G,\bm{h}_{m_{i}}=\phi(m_{i}),\text{ for all $m_{i}$}\in M_{i}.$
Next, an attention operator is used to aggregate the motif encoding using $\bm{h}_{G_{i}}$ as query and $\bm{h}_{m_{i}}$s as keys and values. The resulting encoding, denoted as $\bm{h}^{\prime}_{G_{i}}$, is the aggregated molecular encoding by combining motifs in $M_{i}$.
$e_{G_{i}m_{i}}=g(\bm{W}\bm{h}_{G_{i}},\bm{W}\bm{h}_{m_{i}}),\alpha_{G_{i}m_{i}}=\frac{\exp(e_{G_{i}m_{i}})}{\sum_{m_{i}\in M_{i}}\exp(e_{G_{i}m_{i}})},\bm{h}^{\prime}_{G_{i}}=\sum_{m_{i}\in M_{i}}\alpha_{G_{i}m_{i}}\cdot\bm{h}_{m_{i}},$
where $g(\cdot)$ is an attention operator, and $\bm{W}$ is weight matrix. The target classifier $f(\cdot)$ then uses aggregated encoding to generate predictions. The training process is designed to minimize the cross-entropy loss between the predicted probabilities of the aggregated molecular encoding and those of the original molecular encoding:
$\mathcal{L}=\sum_{i}\mathcal{L}_{CE}(f(\bm{h}_{G_{i}}),f(\bm{h}^{\prime}_{G_{i}})).$ (1)
The attention operator is trained to identify important motifs to represent the molecule. This leads to the molecule-motif relation score $\bm{\alpha}_{G_{i}m_{i}}$ and the probabilities associated with label predictions $\bm{\hat{y}}=f(\bm{h}^{\prime}_{G_{i}})$.
@@
We repeat the process of Stage 1 to compute the molecule-motif scores for each molecular graph. After that, we create a molecule-motif score matrix $\bm{S}^{mm}\in\mathbb{R}^{V\times|\mathcal{G}|}$ and a molecule-class probability matrix $\bm{P}\in\mathbb{R}^{|\mathcal{G}|\times C}$. Here, ${S}^{mm}_{pk}$ represents the molecule-motif relation score between molecule $k$ and motif $p$. If the motif $i$ does not appear in the molecule $j$, we set ${S}^{mm}_{ij}=0$. Element ${P}_{kr}$ denotes the probability that molecule $k$ is associated with label $r$.
Subsequently, we perform a matrix multiplication between $\bm{S}^{mm}$ and $\bm{P}$ to get class-motif relation score matrix $\bm{S}^{cm}$:
$\bm{S}^{cm}=\bm{S}^{mm}\bm{P},{S}^{cm}_{pr}=\sum_{k=1}^{|\mathcal{G}|}{S}^{mm}_{pk}{P}_{kr}$
which represents the relationship between motif $p$ and class $r$. Then, we normalize $\bm{S}^{cm}$ by dividing each column, corresponding to a motif, by the number of molecules it appears.
@@ 3.3 Class-wise Motif-based Graph Generation
Figure 3: Class-wise motif-based graph generation. Starting with a molecular graph, we first construct a junction tree. Next, a tree encoder is applied to obtain a tree encoding, which is then decoded by a tree decoder to reconstruct the junction tree. Finally, a graph decoder uses the predicted junction tree to reproduce the molecular graph.
Section 3.2 identifies motifs of high relevance for each class. Building on this, our methodology employs a motif-based generative model to construct model-level explanations tailored to each class. The motif-based molecular generation method can be divided into VAE-based (Jin et al., 2018; 2020) and RL-based methods (Yang et al., 2021). In our approach, we employ a VAE-based method as a generator. However, our approach can easily adapt to RL-based methods. The graph generation process is illustrated in Figure 3.
@@ 3.3.2 Graph Encoder
We encode the latent representation of $G=(\bm{A},\bm{X})$ by the target GNN feature extractor $\phi(\cdot)$: $\bm{h}_{G}=\phi(\bm{A},\bm{X})$. The mean $\bm{\mu}_{G}$ and log variance $\log\bm{\sigma}_{G}$ of the variational posterior approximation are computed from $\bm{h}_{G}$ with two separate affine layers. $\bm{z}_{G}$ is sampled from a Gaussian distribution $\mathcal{N}(\bm{\mu}_{G},\bm{\sigma}_{G})$. In the variational posterior approximation, we use two different affine layers to compute the mean, represented as $\bm{\mu}_{G}$, and the log variance, denoted as $\log\bm{\sigma}_{G}$, from $\bm{h}_{G}$. Then, we sample $\bm{z}_{G}$ from a Gaussian distribution $\mathcal{N}(\bm{\mu}_{G},\bm{\sigma}_{G})$.